# Teaching Code LLMs to Use Autocompletion Tools in Repository-Level Code Generation

### 12 Jan 2024 v1

@@ Abstract
Recent code large language models (LLMs) have shown promising performance in generating standalone functions but face limitations in repository-level code generation due to their lack of awareness of repository-level dependencies (e.g., user-defined attributes), resulting in dependency errors such as undefined-variable and no-member errors. In this work, we introduce ToolGen, an approach that integrates autocompletion tools into the code LLM generation process to address these dependencies. ToolGen comprises two main phases: Data Augmentation and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). During the offline phase, ToolGen augments functions within a given code corpus with a special mark token, indicating positions to trigger autocompletion tools. These augmented functions, along with their corresponding docstrings, are then used to fine-tune a selected code LLM. In the online phase, ToolGen iteratively generates functions by predicting tokens step-by-step using the fine-tuned LLM. Whenever a mark token is encountered, ToolGen invokes the autocompletion tool to suggest code completions and selects the most appropriate one.
We conduct comprehensive experiments to evaluate ToolGen’s effectiveness in repository-level code generation. To facilitate this evaluation, we create a benchmark comprising 680 real-world code repositories and introduce two new repository-level metrics: Dependency Coverage and Success Rate. The results demonstrate that ToolGen significantly improves dependency coverage by 15.2% to 45.8% and success rates by 10.9% to 42.2% across three distinct code LLMs, while maintaining competitive performance in widely-recognized similarity metrics such as BLEU-4, CodeBLEU, and Edit Similarity. Furthermore, our generalizability evaluation confirms ToolGen’s consistent performance when applied to diverse code LLMs, including various model architectures and scales.
@@ I Introduction
Code generation has been a longstanding focal point in the field of software engineering. Recent advancements have introduced a variety of code large language models (code LLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] constructed upon the Transformer model architecture [14]. These models are either pre-trained or fine-tuned on extensive code corpora, enabling them to automatically generate code based on provided natural language descriptions. These code LLMs have demonstrated notable effectiveness in the generation of code blocks or functions. For instance, CodeLlama [12], built upon the foundational Llama2 model [15], has achieved state-of-the-art results among open code LLMs (e.g., CodeGen [8] and StarCoder [11]), on benchmarks like HumanEval [3] and MBPP [16] that focus on standalone functions.
However, it is crucial to emphasize that in real-world code repositories, more than 70% of functions are not standalone [17]. Code LLMs encounter significant challenges when generating such real-world functions, primarily because they lack an awareness of repository-level dependencies, such as user-defined functions and attributes, during the code generation process [17]. This limitation often leads to the generation of code with dependency errors, including undefined-variable and no-member errors. These errors impede the usability and effectiveness of the code LLMs [18]. For example, consider the scenario depicted in Figure 1. A code LLM (e.g., CodeLlama) might incorrectly predict “_updates” after generating “... self.”, resulting in a no-member error because the object “self” does not possess an attribute named “_updates”.
On the other hand, modern Integrated Development Environments (IDEs) commonly incorporate code autocompletion tools based on program analysis. These tools, like Jedi [19], leverage their ability to analyze the current incomplete function’s state and project context to provide valid completion recommendations. This includes suggestions for accessible variables, attributes, and functions. For instance, when encountering “self.” in Figure 1, Jedi can infer and recommend 68 accessible attributes defined within “self”, including the target suggestion “_registered_updates”. Therefore, if we can seamlessly switch between code LLMs and the use of autocompletion tools, we have the potential to significantly reduce the occurrence of dependency errors in repository-level code generation.
@@
Recent research has delved into the integration of external tools into the generation process of LLMs to mitigate their limitations in constrained generation scenarios. One noteworthy example is ToolFormer [20], which creates an augmented dataset to instruct LLMs on invoking existing arithmetic calculators. This integration effectively reduces errors in generated text involving arithmetic calculations. Building upon ToolFormer’s inspiration, Zhang et al. [21] introduce ToolCoder, an approach designed to teach LLMs how to utilize information-retrieval-based (IR-based) API search tools during the code generation process. While ToolCoder targets the generation of functionally correct standalone functions and demonstrates promising results, the integrated IR-based API search tools do not consider repository-level dependencies, limiting their potential in resolving dependency errors. Another relevant example of harnessing external tools is Repilot [22], which leverages code completion tools to filter out impractical suggestions made by LLMs in the context of automatic program repairing (APR). Unlike repository-level code generation, Repilot’s primary focus is on generating valid bug-fix patches rather than entire functions.
@@
In this work, we aim at integrating program-analysis-based code autocompletion tools into the generation process of code LLMs. Achieving the incorporation presents two key challenges. (i) Determining when to trigger the invocation of autocompletion tools during the generation process: The generation process of LLMs is a step-by-step decoding process where each subsequent token is predicted based on previous tokens. Typically, a function consists of dozens or even hundreds of tokens, making it impractical to invoke code autocompletion tools at every decoding step. In the case of tools like ToolFormer and ToolCoder, ChatGPT is employed to augment the training corpus by introducing special tokens into the text or code to mark positions where tool invocation is needed. After training on this augmented corpus, LLMs can predict the special token at the appropriate step, thereby triggering tool invocation. However, this ChatGPT-based augmentation method is less effective for repository-level code generation due to the presence of repository-level dependencies. The special token must be precisely inserted at positions involving such dependencies, such as when accessing user-defined attributes. (ii) Selecting the target suggestion from the recommended completions of autocompletion tools: Unlike tools like arithmetic calculation or API search integrated into ToolFormer and ToolCoder, which typically return a single result for each invocation, autocompletion tools often provide multiple completion suggestions sorted alphabetically. For instance, as depicted in Figure 1, Jedi returns a list of 68 completion suggestions (excluding builtin attributes), with the target suggestion being the 46th one in the list. Consequently, after invoking autocompletion tools, it is essential to assess the suggestions based on the generated code and select the most appropriate one. Furthermore, this selection process needs to be seamlessly integrated into the code generation process to ensure efficiency and coherence.
@@
To tackle the challenges, we propose ToolGen, an approach to integrate autocompletion tools into the generation process of code LLMs to support repository-level code generation. ToolGen has two main phases: Data Augmentation and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). In the offline phase, ToolGen analyzes source files within a corpus of code repositories, creating abstract syntax trees (ASTs) and extracting function definitions. It augments these functions with a special token, <COMP>, signifying the positions to trigger autocompletion tools. These augmented functions, paired with their respective docstrings, are then employed to fine-tune a selected code LLM. In the online phase, ToolGen iteratively constructs a function based on a provided docstring by predicting tokens step-by-step through the fine-tuned LLM. Whenever a <COMP> token is encountered, ToolGen invokes the autocompletion tool to suggest code completions, drawing from the current repository context. It then selects the most suitable suggestion and appends it to the existing tokens, proceeding with predicting tokens until a termination condition is met.
@@
We conduct extensive experiments to evaluate the effectiveness of ToolGen in repository-level code generation. To facilitate this evaluation, we first construct a benchmark, which includes 12,744 Python functions from 680 real-world code repositories and defines two new repository-level metrics, namely Dependency Coverage and Success Rate. Dependency Coverage quantifies the proportion of repository-level dependencies present in ground-truth functions and successfully covered by the generated functions, while Success Rate measures the percentage of generated functions that pass a dependency error check. he evaluation results demonstrate that ToolGen exhibits comparable performance in widely-recognized similarity metrics such as BLEU-4, CodeBLEU, and Edit Similarity. Importantly, ToolGen achieves significant improvements in dependency coverage, ranging from 15.2% to 45.8%, and success rates, spanning from 10.9% to 42.2%, across three distinct code LLMs. The findings from the generalizability evaluation reaffirm the consistent performance of ToolGen when applied to diverse code LLMs, encompassing various model architectures and scales.
@@
  • An Evaluation Benchmark, which encompasses 12,744 Python functions drawn from 680 real-world code repositories, along with the introduction of two novel repository-level metrics: Dependency Coverage and Success Rate.
  • Experimental Results, which affirm the efficacy of ToolGen in repository-level code generation. ToolGen demonstrates substantial improvements in dependency coverage, ranging from 15.2% to 45.8%, and success rate, spanning from 10.9% to 42.2%, across three distinct code LLMs.
@@ II-A Code LLMs
To ensure that the employed code LLM can recognize and predict the special token <COMP>, we initially incorporate this token into the LLM’s vocabulary, denoted as $\mathbb{V}_{llm}$. This addition results in an expanded vocabulary represented as:
@@ III Approach
We propose an approach named ToolGen to integrate autocompletion tools into the generation process of code LLMs to support repository-level code generation.
@@ III-A Overview
Figure 3 presents an overview of ToolGen, which consists of two main phases, namely (i) Data Augmentation and Model Fine-tuning (Offline) and (ii) Tool-integrated Code Generation (Online).
@@
In data augmentation and model fine-tuning, ToolGen parses each source file in the given code repositories into an abstract syntax tree (AST) and then extract extracts function definitions from the AST; For each extract function definition, ToolGen then utilizes an autocompletion tool to augment it with the special token <COMP> to mark the positions to invoke the tool, and then assembles a pair of docstring and augmented function; After process all code repositories, ToolGen employs the resulting pairs of docstrings and augmented functions to fine-tune a code LLM, resulting in a fine-tuned code LLM that can predict <COMP> at suitable positions to trigger the autocompletion tool.
@@
III-B Code Augmentation and Model Fine-tuning
III-B1 Code Augmentation
We construct an augmented dataset to facilitate the learning process of code LLMs in determining when to utilize autocompletion tools during code generation. In this augmented dataset, the special token <COMP> is inserted at specific locations within code functions, indicating when autocompletion tools should be triggered.
Given a code repository $\mathcal{R}$, we traverse each source file $file$ within it based on the file’s suffix (e.g., .py for Python) and then proceed to analyze the functions defined in the source file. To achieve this, we parse the source file into an abstract syntax tree (AST), where the functions are represented as function-definition nodes. Each function-definition node contains multiple AST-tokens, which are smallest individual units, such as keywords, identifiers, literals, operators, and punctuators, within programming language syntax. These AST-tokens differ from the tokens in the LLM’s vocabulary $\mathbb{V}_{llm}$. Typically, an AST-token comprises one or more tokens from $\mathbb{V}_{llm}$. For example, the AST-token “_registered_updates” consists of six tokens in vocabulary of CodeLlama, i.e., [“_”, “register”, “ed”, “_”, “up”, “dates”].
For each function within the source file, we identify its corresponding function-definition node, denoted as $node$, and apply Algorithm 1 to it. The purpose of this algorithm is to navigate through the function body and identify suitable positions for inserting the special token <COMP>. More specifically, as the algorithm iterates through each AST-token $t$ within the function body $node.body$ (line 2), it performs two crucial checks. First, it employs the isIdentifier procedure to determine whether $t$ is an identifier. Second, it verifies that $t$ is not a built-in attribute, such as “__dict__” in Python, using the isBuiltin procedure. These conditions are essential because dependency errors often arise from user-defined attributes categorized as identifiers rather than other AST-tokens like language keywords. Additionally, these checks prevent the insertion of <COMP> at positions where the code LLM can confidently predict the following tokens, thus minimizing unnecessary tool invocations. When both conditions are met, the algorithm updates the caret position $\mathcal{P}$ to the start position of $t$ (line 4) and invokes the autocompletion tool to obtain a list of completion suggestions, denoted as $\mathbb{C}$ (line 5). If $\mathbb{C}$ contains $t$, indicating that the tool can propose the desired identifier, the special token <COMP> is inserted before $t$ to mark the position for triggering the autocompletion tool (lines 6-7). Upon executing the algorithm, we obtain the augmented function code $\mathcal{F}_{aug}$.
@@ 4             $\mathbb{C}\leftarrow\textsc{tool-complete}(\mathcal{R},\mathcal{P})$
Algorithm 1 Function Code Augmentation
@@ 3             $seq\leftarrow\textsc{llm-tokenize}(comp)$
12             $\bm{p}^{|\mathbb{V}|}\leftarrow\bm{p}^{|\mathbb{V}|}\times\bm{m}^{|\mathbb{V}|}$
@@ 17                         $node\leftarrow child$
Algorithm 3 LLM-based Suggestion Selection
@@ IV-B Approach Implementation
Base Model. In ToolGen, we explore the utilization of three distinct code LLMs to encompass diverse model architectures and parameter scales:
@@
Code Augmentation. To create the augmented dataset for fine-tuning the employed base model, we begin with the Python corpus from the CodeSearchNet dataset. Since the CodeSearchNet dataset does not provide complete code repositories from which to extract Python functions, we initiate the process by crawling the code repositories listed in the dataset. Subsequently, we follow the procedure outlined in Section III-B1 to extract and augment functions within these code repositories, ultimately generating the augmented dataset. It’s important to note that the CodeSearchNet dataset includes a partitioning into training, validation, and test sets. For our code augmentation process, we exclusively utilize the code repositories associated with the training set. The resulting augmented dataset comprises a total of 256,870 pairs of docstrings and augmented functions, which are sourced from 12,231 distinct Python code repositories. Regarding dataset statistics, the average token count for docstrings is 10.96, and for augmented functions, it is 54.13. Additionally, the special token <COMP> appears an average of 5.37 times within these functions.

### 21 Jan 2024 v2

@@ Abstract
Recent code large language models (LLMs) have shown promising performance in generating standalone functions. However, they face limitations in repository-level code generation due to their lack of awareness of repository-level dependencies (e.g., user-defined attributes), resulting in dependency errors such as undefined-variable and no-member errors. In this work, we introduce ToolGen, an approach that integrates autocompletion tools into the code LLM generation process to address these dependencies. ToolGen comprises two main phases: Trigger Insertion and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). During the offline phase, ToolGen augments functions within a given code corpus with a special mark token, indicating positions to trigger autocompletion tools. These augmented functions, along with their corresponding docstrings, are then used to fine-tune a selected code LLM. In the online phase, ToolGen iteratively generates functions by predicting tokens step-by-step using the fine-tuned LLM. Whenever a mark token is encountered, ToolGen invokes the autocompletion tool to suggest code completions and selects the most appropriate one through constrained greedy search.
We conduct comprehensive experiments to evaluate ToolGen’s effectiveness in repository-level code generation. To facilitate this evaluation, we create a benchmark comprising 680 real-world code repositories and introduce two new repository-level metrics: Dependency Coverage and Static Validity Rate. The results demonstrate that ToolGen significantly improves Dependency Coverage by 15.2% to 45.8% and Static Validity Rate by 10.9% to 42.2% across three distinct code LLMs, while maintaining competitive performance in widely-recognized similarity metrics such as BLEU-4, CodeBLEU, and Edit Similarity. Furthermore, our generalizability evaluation confirms ToolGen’s consistent performance when applied to diverse code LLMs, including various model architectures and scales.
@@ I Introduction
Code generation has been a longstanding focal point in the field of software engineering. Recent advancements have introduced a variety of code large language models (LLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] constructed upon the Transformer model architecture [14]. These models are either pre-trained or fine-tuned on extensive code corpora, enabling them to automatically generate code based on provided natural language descriptions. These code LLMs have demonstrated notable effectiveness in the generation of code blocks or functions. For instance, CodeLlama [12], built upon the foundational Llama2 model [15], has achieved state-of-the-art results among open code LLMs (e.g., CodeGen [8] and StarCoder [11]), on benchmarks like HumanEval [3] and MBPP [16] that focus on standalone functions.
However, it is crucial to emphasize that in real-world code repositories, more than 70% of functions are not standalone [17]. Code LLMs encounter significant challenges when generating such real-world functions, primarily because they cannot be aware of repository-level dependencies, such as user-defined functions and attributes, during the code generation process [17]. This limitation often leads to the generation of code with dependency errors, including undefined-variable and no-member errors. These errors impede the usability and effectiveness of the code LLMs [18]. For example, consider the scenario depicted in Figure 1. A code LLM (e.g., CodeLlama) might incorrectly predict “_updates” after generating “... self.”, resulting in a no-member error because the object “self” does not possess an attribute named “_updates”.
Meanwhile, modern Integrated Development Environments (IDEs) take a different approach, which typically incorporates code autocompletion tools based on program analysis. These tools, like Jedi [19], leverage their ability to analyze the current incomplete function’s state and project context to provide valid completion recommendations. This includes suggestions for accessible variables, attributes, and functions. For instance, when encountering “self.” in Figure 1, Jedi can infer and recommend 68 accessible attributes defined within “self”, including the target suggestion “_registered_updates”. Therefore, if we can seamlessly switch between code LLMs and the use of autocompletion tools, we have the potential to significantly reduce the occurrence of dependency errors in repository-level code generation.
@@
In fact, recent research has delved into the integration of external tools into the generation process of LLMs to mitigate their limitations in constrained generation scenarios. One noteworthy example is ToolFormer [20], which creates an augmented dataset to instruct LLMs on invoking existing arithmetic calculators. This integration effectively reduces errors in generated text involving arithmetic calculations. Building upon ToolFormer’s inspiration, Zhang et al. [21] introduce ToolCoder, an approach designed to teach LLMs how to utilize information-retrieval-based (IR-based) API search tools during the code generation process. While ToolCoder targets the generation of functionally correct standalone functions and demonstrates promising results, the integrated IR-based API search tools do not consider repository-level dependencies, limiting their potential in resolving dependency errors. Another relevant example of harnessing external tools is Repilot [22], which leverages code completion tools to filter out impractical suggestions made by LLMs in the context of automatic program repairing (APR). Unlike repository-level code generation, Repilot’s primary focus is on generating valid bug-fix patches rather than entire functions.
@@
In this work, we aim at integrating program-analysis-based code autocompletion tools into the generation process of code LLMs. Achieving the incorporation presents two key challenges. (i) Determining when to trigger the invocation of autocompletion tools during the generation process: The generation process of LLMs is a step-by-step decoding process where each subsequent token is predicted based on previous tokens. In general, a function consists of dozens or even hundreds of tokens, making it impractical to invoke code autocompletion tools at every decoding step. In the case of tools like ToolFormer and ToolCoder, ChatGPT is employed to augment the training corpus by introducing special tokens into the text or code to mark positions where tool invocation is needed. After training on this augmented corpus, LLMs can predict the special token at the appropriate step, thereby triggering tool invocation. However, this ChatGPT-based augmentation method is less effective for repository-level code generation due to the presence of repository-level dependencies. The special token must be precisely inserted at positions involving such dependencies, such as when accessing user-defined variables. (ii) Selecting the target suggestion from the recommended completions of autocompletion tools: Different from tools like arithmetic calculation or API search integrated into ToolFormer and ToolCoder, which return a single result for each invocation, autocompletion tools often provide multiple completion suggestions sorted alphabetically. For instance, as depicted in Figure 1, Jedi returns a list of 68 completion suggestions (excluding builtin attributes), with the target suggestion being the 46th one in the list. Consequently, after invoking autocompletion tools, it is essential to assess the suggestions based on the generated code and select the most appropriate one. Furthermore, this selection process needs to be seamlessly integrated into the code generation process to ensure efficiency and coherence.
@@
To tackle the challenges, we propose ToolGen, an approach to integrate autocompletion tools into the generation process of code LLMs to support repository-level code generation. ToolGen has two main phases: Trigger Insertion and Model Fine-tuning (Offline), and Tool-integrated Code Generation (Online). In the offline phase, ToolGen analyzes source files within a corpus of code repositories, creating abstract syntax trees (ASTs) and extracting function definitions. It augments these functions by inserting a special token, <COMP>, signifying the positions to trigger autocompletion tools. The insertion positions are established by navigating through the functions and identifying the identifiers that can be recommended by autocomplete tools. These augmented functions, paired with their respective docstrings, are then employed to fine-tune a selected code LLM. In the online phase, ToolGen iteratively constructs a function based on a provided docstring by predicting tokens step-by-step through the fine-tuned LLM. Whenever a <COMP> token is encountered, ToolGen invokes the autocompletion tool to suggest code completions, drawing from the current repository context. Subsequently, it identifies the most appropriate suggestion through a constraint greedy search algorithm, appending this selected suggestion to the current tokens. This process continues as it predicts tokens until a specified termination condition is satisfied.
@@
We conduct extensive experiments to evaluate the effectiveness of ToolGen in repository-level code generation. To facilitate this evaluation, we first construct a benchmark, which includes 12,744 Python functions from 680 real-world code repositories and defines two new repository-level metrics, namely Dependency Coverage and Static Validity Rate. Dependency Coverage quantifies the proportion of repository-level dependencies present in ground-truth functions and successfully covered by the generated functions, while Static Validity Rate measures the percentage of generated functions that pass a dependency error check. The evaluation results demonstrate that ToolGen exhibits comparable performance in widely-recognized similarity metrics such as BLEU-4, CodeBLEU, and Edit Similarity. Importantly, ToolGen achieves significant improvements in Dependency Coverage, ranging from 15.2% to 45.8%, and Static Validity Rate, spanning from 10.9% to 42.2%, across three distinct code LLMs. Moreover, the results from our generalizability evaluation confirm that ToolGen consistently performs well across a variety of code LLMs, with different model architectures and scales.
@@
  • An Evaluation Benchmark, which encompasses 12,744 Python functions drawn from 680 real-world code repositories, along with the introduction of two novel repository-level metrics: Dependency Coverage and Static Validity Rate.
  • Extensive Experimental Results, which affirm the efficacy of ToolGen in repository-level code generation. ToolGen demonstrates substantial improvements in Dependency Coverage, ranging from 15.2% to 45.8%, and Static Validity Rate, spanning from 10.9% to 42.2%, across three distinct code LLMs.
@@ II-A Code LLMs
On top of the standard generation process, to ensure that the employed code LLM can recognize and predict the special token <COMP>, we initially incorporate this token into the LLM’s vocabulary, denoted as $\mathbb{V}_{llm}$. Formally, this addition results in an expanded vocabulary represented as:
@@ III Approach
In this section, we elaborate on our approach named ToolGen to integrate autocompletion tools into the generation process of code LLMs to support repository-level code generation.
@@ III-A Overview
Figure 3 presents an overview of ToolGen, which consists of two main phases, namely (i) Trigger Insertion and Model Fine-tuning (Offline) and (ii) Tool-integrated Code Generation (Online).
@@
In trigger insertion and model fine-tuning, ToolGen parses each source file in the given code repositories into an abstract syntax tree (AST) and then extract extracts function definitions from the AST; For each extract function definition, ToolGen then utilizes an autocompletion tool to augment it with the special token <COMP> to mark the positions to invoke the tool, and then assembles a pair of docstring and augmented function; After process all code repositories, ToolGen employs the resulting pairs of docstrings and augmented functions to fine-tune a code LLM, resulting in a fine-tuned code LLM that can predict <COMP> at suitable positions to trigger the autocompletion tool.
@@
III-B Trigger Insertion and Model Fine-tuning
III-B1 Trigger Insertion
We employ a trigger insertion method to facilitate the learning process of code LLMs in determining when to utilize autocompletion tools during code generation. In this method, the special token <COMP> is inserted at specific locations within code functions, indicating when autocompletion tools should be triggered.
Given a code repository $\mathcal{R}$, we traverse each source file $file$ within it based on the file’s suffix (e.g., .py for Python) and then proceed to analyze the functions defined in the source file. To achieve this, we parse the source file into an abstract syntax tree (AST), where the functions are represented as function-definition nodes. Each function-definition node contains multiple AST-tokens, which are smallest individual units, such as keywords, identifiers, literals, operators, and punctuators, within programming language syntax. Note that these AST-tokens differ from the tokens in the LLM’s vocabulary $\mathbb{V}_{llm}$. Typically, an AST-token comprises one or more tokens from $\mathbb{V}_{llm}$. For example, the AST-token “_registered_updates” consists of six tokens in vocabulary of CodeLlama, i.e., [“_”, “register”, “ed”, “_”, “up”, “dates”].
For each function within the source file, we identify its corresponding function-definition node, denoted as $node$, and apply Algorithm 1 to it. The purpose of this algorithm is to traverse the function body and identify specific identifiers that are eligible for suggestions by autocomplete tools. Subsequently, the special token <COMP> is inserted in front of these chosen identifiers. More specifically, as the algorithm iterates through each AST-token $t$ within the function body $node.body$ (line 2), it performs two crucial checks. First, it employs the isIdentifier procedure to determine whether $t$ is an identifier. Second, it verifies that $t$ is not a built-in attribute, such as “__dict__” in Python, using the isBuiltin procedure. These conditions are essential because dependency errors often arise from user-defined attributes categorized as identifiers rather than other AST-tokens like language keywords. Additionally, these checks prevent the insertion of <COMP> at positions where the code LLM can confidently predict the following tokens, thus minimizing unnecessary tool invocations. When both conditions are met, the algorithm updates the caret position $\mathcal{P}$ to the start position of $t$ (line 4) and invokes the autocompletion tool to obtain a list of completion suggestions, denoted as $\mathbb{C}$ (line 5). If $\mathbb{C}$ contains $t$, indicating that the tool can propose the desired identifier, the special token <COMP> is inserted before $t$ to mark the position for triggering the autocompletion tool (lines 6-7). Upon executing the algorithm, we obtain the augmented function code $\mathcal{F}_{aug}$.
@@ 4             $\mathbb{C}\leftarrow\textsc{tool-complete}(\mathcal{R},\mathcal{P})$
Algorithm 1 Trigger Insertion
@@ 3             $seq\leftarrow\textsc{llm-tokenize}(comp)$
12             $\bm{p}^{|\mathbb{V}|}\leftarrow\bm{p}^{|\mathbb{V}|}\odot\bm{m}^{|\mathbb{V}|}$
@@ 17                         $node\leftarrow child$
Algorithm 3 Suggestion Selection based on Constraint Greedy Search
@@ IV-B Implementation
Base Model. In ToolGen, we explore the utilization of three distinct code LLMs to encompass diverse model architectures and parameter scales. These code LLMs demonstrate impressive performance in code generation and have found extensive utilization in prior studies [23, Compilable, 4, 5, 12] for fine-tuning and evaluation.
@@
Trigger Insertion. To create the augmented dataset for fine-tuning the employed base model, we begin with the Python corpus from the CodeSearchNet dataset. Since the CodeSearchNet dataset does not provide complete code repositories from which to extract Python functions, we initiate the process by crawling the code repositories listed in the dataset. Subsequently, we follow the procedure outlined in Section III-B1 to extract and augment functions within these code repositories, ultimately generating the augmented dataset. It’s important to note that the CodeSearchNet dataset includes a partitioning into training, validation, and test sets. For our trigger insertion process, we exclusively utilize the code repositories associated with the training set. The resulting augmented dataset comprises a total of 256,870 pairs of docstrings and augmented functions, which are sourced from 12,231 distinct Python code repositories. Regarding dataset statistics, the average token count for docstrings is 10.96, and for augmented functions, it is 54.13. Additionally, the special token <COMP> appears an average of 5.37 times within these functions.